OpenStudy (anonymous):

find an equation for the line with the given properties Question: Perpendicular to the line x − 2y = −5; containing the point (0, 4)

zepdrix (zepdrix):

\(\Large \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

Ok tortilla! Let's see what we've got here. Let's take the line they gave us, and put it into a form that's easier to read.\[\Large x-2y=-5\]Subtract x from each side,\[\Large -2y=-x-5\]Multiply both sides by -1,\[\Large 2y=x+5\]Divide by 2,\[\Large y=\frac{1}{2}x+\frac{5}{2}\]

zepdrix (zepdrix):

From here, we can see that this line has a slope of 1/2. For a line to be perpendicular to this one, it's slope will relate to this slope in a specific way. Do you remember how? :O